OpenStudy (anonymous):

Using DeMoivre's Theorem, calculate ( sqrt(3) + i )^4. I'm aware DeMoivre's Thm: z^n=r^n (cos(ntheta)+isin(ntheta))

OpenStudy (anonymous):

i have r=2 that's it lol

OpenStudy (anonymous):

i have the triangle but not sure how to move on

OpenStudy (kinggeorge):

You don't have that much further to go. Since you have the triangle, you should have something like this.All you need to do, is find the value of x, where x is that angle.

OpenStudy (anonymous):

i can't use a calculator

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

cos(theta)=sqrt(3)/2

OpenStudy (anonymous):

theta=30 degrees

OpenStudy (anonymous):

is that correct?

OpenStudy (kinggeorge):

I believe that would be correct.

OpenStudy (kinggeorge):

Unless satellite disagrees? Have I been too hasty?

OpenStudy (anonymous):

thanks sorry i usually don't have so many questions too much stress for tomorrows Math CSET 112 :-(

OpenStudy (anonymous):

just to butt in for a second, since this is an adult problem, it really should be \[\theta =\frac{\pi}{6}\] but i guess you can use 30

OpenStudy (kinggeorge):

Good point. You should usually have it as \(\pi/6\) when using cosines and sines.